OpenStudy (anonymous):

You are knitting a blanket. You want the area of the blanket to be 24 ft^2. You want the length of the blanket to be 2 ft longer than its width. What should the dimensions of the blanket be?

OpenStudy (anonymous):

2W=L L*W=24

OpenStudy (anonymous):

Ok,so itd be (2w)w=24?

OpenStudy (anonymous):

It's two feet longer. Not two times longer

OpenStudy (anonymous):

dang, so it would be W+2=L

OpenStudy (anonymous):

so you would then have (W+2)(W)=24

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

no prob :) and thanks @Xavier for catching my mistake!